Find at least one solution to the following equation:
sin(x^2 − 1)/1 − sin(x^2 − 1) = sin(x) + sin^2(x) + sin^3(x) + sin^4(x) + · · ·
1
Expert's answer
2020-11-17T06:08:38-0500

Right side of the equation is a geometric series with "a = sin(x)" and "r = \frac{sin^2(x)}{sin(x)} = sin(x)".

The sum of a geometric series:


"S = \frac{a}{1-r}""S = \frac{sin(x)}{1-sin(x)}"

Hence the equation is:


"\frac{sin(x^2-1)}{1-sin(x^2-1)} = \frac{sin(x)}{1-sin(x)}"

"sin(x^2-1)(1-sin(x)) = sin(x)(1-sin(x^2-1))"

"sin(x^2-1) = sin(x)"

Since the period the function "y = sin(x)" is "2\pi", the equation is:



"x^2-1 = x +2\pi n"

"x^2-x-(1+2\pi n) = 0"

"x_1 = \frac{1-\sqrt{5+8\pi n }}{2}, x_2 = \frac{1+\sqrt{5+8\pi n }}{2}, n = 0, 1, 2, ..."


In particular, for n = 0 we have:


"x_1 = \frac{1-\sqrt{5}}{2}, x_2 = \frac{1+\sqrt{5}}{2}"

For n = 1:


"x_1 = \frac{1-\sqrt{5+8\pi}}{2}, x_2 = \frac{1+\sqrt{5+8\pi}}{2}"